Bacterial cells grow at a rate proportional to the volume of dividing cells at that moment. If V0

is the volume of dividing cells at any time t . Find the time at which the volume of the cells will
be double its original size.
1
Expert's answer
2020-05-21T12:23:13-0400

"\frac{dV(t)}{dt}=V(t)" then "\ln" |"V"|"=t+\ln C" then "V=Ce^t" . As "V(0)=V_0=C," then

"V(t)=V_0e^t" hence "V(\tau)=2V_0=V_0e^\tau" hence "e^\tau=2," hence "\tau=
ln 2\approx0.69 s"